OpenStudy (anonymous):

Are the graphs of the lines in the pair parallel? Explain. y = 1/6 x 1+ 8 –2x + 12y = –11

OpenStudy (hoblos):

–2x + 12y = –11 12y = 2x-11 y= 1/6 x - 11/12 so for both lines the slope = 1/6 so they are parallel

OpenStudy (anonymous):

Yes, since the slopes are the same and the y-intercepts are the same. or Yes, since the slopes are the same and the y-intercepts are different.

OpenStudy (hoblos):

the y-intercepts are different... anyway the y-intercept dont affect parallelism ;)